OpenStudy (anonymous):

Which of the following represents a perfect square trinomial? 9x2 + 24x + 16 x2 + 8x – 16 x2 + 8x + 32 x2 + 8x + 36

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Do you know factoring, Lou?

OpenStudy (anonymous):

Yeah.

Parth (parthkohli):

Can you factor all the four expressions?

OpenStudy (anonymous):

Probably yeah.

Parth (parthkohli):

Can you factor all and list them down? :)

OpenStudy (anonymous):

K give me a sec. :)

OpenStudy (anonymous):

Hey wait, for factoring. It's 2 numbers that when multiplied equal the last number, and add up to the middle one right?

Parth (parthkohli):

I'd give you example for the fourth one. x^2 + 8x + 36 We need to find two numbers that add up to 8 and multiply to get 36. 6 * 6 = 36 4 * 9 = 36 12 * 3 = 36 But there are no numbers we can find so x^2 + 8x + 36 is prime.

OpenStudy (anonymous):

oh okay yeah. Lol was spooked for a sec. wasn't thinking about prime's. :)

Parth (parthkohli):

Another way to determine these is to find the discriminant.

Parth (parthkohli):

Leave that method.

OpenStudy (anonymous):

K, I'm thinking they're all prime.

Parth (parthkohli):

Let's try to factor 9x^2 + 24x + 16 16 * 9 = 144 We will now try to find two numbers that multiply to get 144 and add to get 24. 12 * 12 = 144 12 + 12 = 24 So now we split the middle term. 9x^2 + 12x + 12x + 16 Factor out 3x from first two and last two. 3x(3x + 4) + 4(3x + 4) Group them to finish with (3x + 4)(3x + 4)

OpenStudy (anonymous):

K. and that would be the answer cause it's (3x+4)^2

Parth (parthkohli):

(3x + 4)(3x + 4) is the same as (3x + 4) x (3x + 4) and when you multiply two numbers that are the same then you can write it as a perfect square.

Parth (parthkohli):

Yep! :)

OpenStudy (anonymous):

Tyvm @ParthKohli and sry if it took a while :)

Parth (parthkohli):

Haha lol no